RHIC Startup: Reducing the beam rigidity with F6 did not
allow for stable running. So it was decided that there is no acceptable
alternative to going in to the A5 to shorten the pulse. It may be possible
to change the tap on the A5 PFN to shorten the pulse. Probably only the
central module of the A5 will need to be pulled out, allowing access to the
other two modules that way.
So the schedule for 9 Nov is:
0400:
Beam activity ends, CA LOTO to be applied to AGS.
0410: Close AGS Vacuum (sector) Valves at A3 & A13 (pet)
0500:
CA Access to AGS for HP survey of A5 kicker.
0530 STAR and Phenix to RA before 0600 -- get Phenix plug door opened.
0600:
Pulsed Power Supply Group (A5 kicker) personnel to enter ring.
0800 RHIC PASS zone 2 (1z1, 1z2, 2z2) to RA at 0800 or earlier
Before1000:
All personnel out of ring -- REVIEW GATE LOG SHEET TO MAKE SURE
ALL ARE ACCOUNTED FOR.
1000-1200:
re-access ring as necessary.
1200:
Remove LOTO from AGS.
1300: Restore beam to W dump.
It is anticipated that the A% work can be done in parallel to the Yellow bus work.
Last night there was an accidental quench of the DX at 5022A. After the quench, there were several cycles of the DX shutoff at 4146A.
AGS RF cavities kept tripping off because the RF was left in CW mode. This indicates that the RF needs to be included in the sequencer.
AtR BPMs are not triggering properly. Loss Monitors do not work in the first turn mode. They need circulating beam to work.
As of 8 Nov 6:45:45 there was circulating deuterons in Blue with good lifetime. Intensities of 0.5E11 were recorded with 85% injection efficiency. RF was turned on at 7:51:30. We need the Abort Kicker and Loss Monitors working for tonight.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email:sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |